Staff:

LBTO: David G, Alex B, Dave T, Juan Carlos G.

Observer: Enrico P, Fabio R, Guido A @ Arcetri (remote)

Forecast.

forecast1_20230220.png

forecast0_20230220.png

Details.

Remain closed. Foggy. bad forecast.

10:59PM

Still closed.

all_sky_20230220_0.png

Power on WFS 11:30PM
  • Bayside stage timeout.
  • Power cycle 24V
  • OK
Chamber temperautre 4C

adsec CP:2.8

Adsec REf Body:3.5

OCAM

two noisy quadrants.
  • Power cycle BCU2K
  • Reconfigure bin.
Adsec flat.

06:59:41.629126| MAIN > Maximum mode index used to apply flat -> # 623

12:01 am

test presetAO.
  • mag 10
  • dist s1.5@1700Hz.
  • 07:03:12.739808| MAIN > Timeout error on baysidex (-5001) TIMEOUT_ERROR.
  • The bayside stage stopped at half way of the movements.
  • Home.
12:06am

PresetAO.
  • mag 10
  • dist s1.5@1700Hz.
  • TN:20230221_072653,_072742,_072822,_072852
12:30am

Enrico joins.

Setup the AO table for the default
  • table_LUCIFER_ACE-AO.txt -> table_LUCIFER_ACE-AO_20230130_500m_v01.txt

Bin4 delay.

1.0 check the bin4.

Bin3 faint end performance

Bin2 mag 13

PresetAO

mag 13.5

shutter:90

optimize:20230221_075044.

Gopt failed.

Process auxloop didn;t start. It reports disk full.

Check disc wfs.

[AOeng@soul-dxwfs source]$ df -h
Filesystem Size Used Avail Use% Mounted on
/dev/mapper/vg00-root
5.8G 5.1G 489M 92% /
tmpfs 7.8G 0 7.8G 0% /dev/shm
/dev/sda1 190M 39M 141M 22% /boot
/dev/mapper/vg00-home
815G 774G 0 100% /home
/dev/mapper/vg00-tmp 976M 6.8M 918M 1% /tmp
/dev/mapper/vg00-var 3.9G 442M 3.2G 12% /var
192.168.39.55:/local/aomeas
7.3T 5.7T 1.2T 84% /local/aomeas
192.168.39.20:/volume2/lbto_flao
49T 28T 22T 57% /ao-data
[AOeng@soul-dxwfs source]$

move the local/aolog to the ao-data/Log_and_Telemetry

Recomended to have a cron to check the disk space of the WFS machines.

du -sh * | sort -hr | head -n10
[AOeng@soul-dxwfs bin]$ du -sh * | sort -hr | head -n10
573M    ModalplotTask.log
553M    core.9996
553M    core.9412
553M    core.8414
553M    core.83791
553M    core.82153
553M    core.80363
553M    core.78655
553M    core.77005
553M    core.75311

1:13am Fabio connected, checking the files.

clean the core dumps.

[AOeng@soul-dxwfs bin]$ df -h
Filesystem Size Used Avail Use% Mounted on
/dev/mapper/vg00-root
5.8G 5.1G 489M 92% /
tmpfs 7.8G 0 7.8G 0% /dev/shm
/dev/sda1 190M 39M 141M 22% /boot
/dev/mapper/vg00-home
815G 333G 441G 44% /home
/dev/mapper/vg00-tmp 976M 6.8M 918M 1% /tmp
/dev/mapper/vg00-var 3.9G 442M 3.2G 12% /var
192.168.39.55:/local/aomeas
7.3T 5.7T 1.2T 84% /local/aomeas
192.168.39.20:/volume2/lbto_flao
49T 28T 22T 57% /ao-data

Open an issue track.

Issue track: #8833

PresetAO

the auxloop doesn;t update.

Restart the WFS soft

PresetaO
  • mag preset13.5
  • shutter 95
  • mag
  • dis s1.5@900
  • optimize: 20230221_08
  • TN:20230221_084210,_084230,_084250,_084311,_084325,_084344.
  • modalSR 60-61
1:46am

presetao
  • mag perset 13.5
  • s2.0@900Hz.
  • Optimize:_085031
  • TN:_085251,_085308,_085345,_085413,_085428,_085510,_085533
  • modalSR:48
  • skip frames ~5%
  • optimize:085620
  • 3.57,0.74,0.48
  • TN:_085828,_085846,_085902,_085924,_085952,_090013
  • vibration at 13Hz
2:03am

Bin4 delay

presetaO
  • mag preset 16.51
  • dist s1.5@500
  • shutter 95 keeping bright
  • rise manual gains.
  • TT:2.5 oscillation at 37Hz
  • TN:091323,091336,_091348
  • working in a new pupil wth different parameters to explore the delay of the slope computer computations.
02:48am

found a combination where the TT gain rised up to 5 and the osclallation frequency is around 100Hz, the frequency expected.

Fabio has to documented the changes how to propagate the change.

03:20am

PresetAO
  • preset mag 16.51
  • bin4 shutter 95
  • mag 13.9
  • optimize gain.
  • observed oscillation at 87Hz for the tip.
  • change the filter to pure integrator and it observed the same freq.
03:50 am

new table to test.

table_LUCIFER_ACE-AO.txt -> table_LUCIFER_ACE-AO_20230221_500m_TESTING01.txt

presetao
  • preset mag 16.51
  • bin4 shutter 95
  • mag 13.9
  • bin4, 1000Hz.
  • optimize:105804
  • TT 1.48 small osc 87Hz
  • TT 2.54 osc 127Hz.
  • increasing the TT gain the 127Hz is excited, visible the two axis oscillating.

04:22am

PresetAO
  • preset mag16.51
  • shutter 95
  • mag 13.9
  • bin4, 500Hz
  • manual optimize
04:32am

PresetAO
  • preset mag16.51
  • shutter 95
  • mag 13.9
  • bin4, 750Hz
  • optimize:_113403
  • slope zero
  • ?
slope_zero-20230221.png

04:38am

PresetAO
  • preset mag16.51
  • shutter 95
  • mag 13.9
  • bin4, 750Hz
  • manual gain TT 3.73 to oscillate, freq 125Hz
04:48am

presetAO
  • preset mag16.51
  • shutter 80
  • mag 16.6
  • Bin4 645Hz
  • s1.5@750Hz
  • cancel preset. due to a nlow or nothing correcting
04:56am
  • preset mag16.51
  • shutter 80
  • s1.5@750Hz
  • mag 16.6
  • TN:_120158,_120213,_120229

Stop 5:00 am

-- JuanCarlosGuerra - 21 Feb 2023
Topic revision: r4 - 21 Feb 2023, JuanCarlosGuerra
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback